OpenStudy (anonymous):

The number of eggs, X, laid by the female tawny owl has the probability distribution x = 2,3,4. P(X=x) = 0.1, 0.2, 0.3. For any egg the probability that it is hatched is 0.8, independently of all other eggs. Obtain th probability distribution of Y, the number of hatched eggs in arandomly chosen nest. Obtain the probability distribution of Y

OpenStudy (kropot72):

\[P(y)=P(x)\times P(hatched)\] \[P(y)=P(x)\times 0.8\] Now you can calculate P(y) for each value of x to find the probability distribution of Y.